# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation with application to sigmoidal classification models

### v2

@@ Abstract
We introduce gradient-flow-guided adaptive importance sampling (IS) transformations for stabilizing Monte-Carlo approximations of leave-one-out (LOO) cross-validated predictions for Bayesian models. After defining two variational problems, we derive corresponding simple nonlinear transformations that utilize gradient information to shift a model’s pre-trained full-data posterior closer to the target LOO posterior predictive distributions. In doing so, the transformations stabilize importance weights. The resulting Monte Carlo integrals depend on Jacobian determinants with respect to the model Hessian. We derive closed-form exact formulae for these Jacobian determinants in the cases of logistic regression and shallow ReLU-activated artificial neural networks, and provide a simple approximation that sidesteps the need to compute full Hessian matrices and their spectra. We test the methodology on an $n\ll p$ dataset that is known to produce unstable LOO IS weights.
@@ 1 Introduction
In Bayesian inference, multiple models are often fit to data and then a selection procedure is applied to decide which model is best. A commonly used criterion for selection is how well a model accurately predicts out-of-sample data, which automatically optimizes fit while avoiding over-fitting. Prediction accuracy is most naturally estimated using cross-validation (CV). A model is trained with some of the data and tested against the unused data. However, estimates of out-of-sample model metrics using train-test splitting are noisy [Dietterich, 1998, Kohavi, 1995] unless computationally expensive k-fold cross-validation (i.e. fitting the model multiple times on cross-over the entire dataset) is employed [Rodriguez et al., 2010, Wong and Yeh, 2020].
Although $N$-fold, also known as leave one out (LOO), CV is the most expensive, there exist computationally efficient LOO techniques that completely avoid refitting. For example, the Akaike Information criteria (AIC) and Bayesian variants [Stone, 1977, Watanabe, 2010, Gelman et al., 2014, Watanabe, 2013] are asymptotic approximations of LOO CV. For Bayesian models, a more precise way to perform LOO CV is to use importance sampling [Vehtari et al., 2017, Piironen and Vehtari, 2017a], which works by using the full data posterior measure as a proposal distribution for each data point’s LOO posterior measure. These methods preclude refitting the model by approximating each LOO posterior measure using full data posterior measure. However, in cases where the LOO measure and full measure are very different, importance sampling can fail [Piironen and Vehtari, 2017a]. To ameliorate this possibility, we introduce an adaptive importance sampling method for LOO CV based on using transformations that bring the proposal distribution closer to LOO posteriors. We derive these transformations by defining gradient flows that minimize given statistical objective. While the transformations are model-dependent, the method is made completely general when using autograd for computing model gradients – for computational efficiency we derive the transformations exactly for a large class of classification models.
@@ 2 Preliminaries
Notation: We denote vectors (assumed to be column vectors unless otherwise stated) using bold-faced lowercase symbols, and matrices using bold-faced uppercase symbols. Given a matrix $\mathbf{W}=(w_{ij})$, the $i$-th row is denoted $\mathbf{w}_{i,}$ and $j$-th column is denoted $\mathbf{w}_{,j}$.
We refer to the entire set of observed training data as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. As shorthand, we denote the set of training data where the i-th observation is left out as $\mathcal{D}^{(-i)}=\mathcal{D}\setminus\{(\mathbf{x}_{i},y_{i})\}$. Expectations with respect to the posterior distribution of $\boldsymbol{\theta}$ are denoted $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}$, and with respect to the posterior distribution of $\boldsymbol{\theta}$ if observation $i$ is left out are denoted $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}$.
@@
Importance sampling-based approximate leave one out cross validation (IS-LOO): Suppose that one has pre-trained a Bayesian model such that one is able to sample its posterior parameters
@@
Figure 1: Relationships between probability densities. One wants to sample from $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)}),$ the LOO distribution for observation $i$, by sampling from the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D}).$ The transformation $T_{i}$ on the full-data posterior brings the sampling distribution closer to the target LOO distribution.
@@
$\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D}).$ Our objective is to use knowledge of this full-data posterior distribution to estimate how the model would behave if any single point is left out at training. One can relate the full-data model to the model with observation $i$ left out using the Bayesian update equation
$\pi(\boldsymbol{\theta}|\mathcal{D})=\frac{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\int\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\boldsymbol{\theta}},$ (1)
@@
which is a Fredholm integral equation of thge with respect second-kind to $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$. This integral equation is in-practice difficult to solve due to the typically-high dimensionality of $\boldsymbol{\theta}.$
@@
Rather than directly inverting Eq. 1, our starting point is the observation that Eq. 1 implies
$\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}=\frac{\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})]}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\equiv\nu_{i}(\boldsymbol{\theta}),$ (2)
@@
which provides the ratio of densities between a distribution we know (the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$) and a distribution whose statistics we would like to compute (the point-wise LOO posterior $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$). To use the former to compute statistics of the latter we turn to Monte-Carlo [Barbu and Zhu, 2020, Robert and Casella, 2013] - the use of statistical sampling to compute a desired quantity (typically an integral). Importance Sampling (IS) is a Monte-Carlo method where one computes expectations with respect to a target distribution by taking a weighted average of samples with respect to a given proposal distribution. For an integrable function $f$,
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}\left[f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\right].$ (3)
@@
We approximate Eq. 3 by sampling over $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, and computing the Monte-Carlo integral
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\approx\sum_{k=1}^{s}\nu_{ik}f(\boldsymbol{\theta}_{k})$ (4)
@@
$\nu_{ik}=\frac{\nu_{i}(\boldsymbol{\theta}_{k})}{\sum_{j=1}^{s}\nu_{i}(\boldsymbol{\theta}_{j})}=\frac{(\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}))^{-1}}{\sum_{k=1}^{s}(\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}))^{-1}},$ (5)
so that the undetermined constant $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\right]$ cancels out. Eqs. 45 define a well-known [Gelfand et al., 1992] Monte-Carlo estimator for LOO.
LOO-based metrics: A Bayesian analogue to the Aikaike Information Criterion (AIC) can be computed using the LOO information criterion (LOO-IC):
LOO-IC $=-2\sum_{i=1}^{n}\log\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left(\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\right)$
$\approx-2\sum_{i=1}^{n}\log\sum_{k=1}^{s}{\nu}_{ik}\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}).$ (6)
Alternatively, for classification problems, one often desires other metrics such as an estimate of the out-of-sample area under the receiver operator curve or precision-recall curve. To compute these quantities, one simply propagates LOO estimates of the outcome probabilities
$\widehat{\mathscr{p}}_{\textrm{loo},i}=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[\mathscr{p}_{i}(\boldsymbol{\theta})\right]\approx\sum_{k=1}^{s}{\nu}_{ik}\mathscr{p}(\boldsymbol{\theta}_{k},\mathbf{x}_{i}),$ (7)
into the relevant formulae. For instance, one may use the probabilistic interpretation of the AUROC [Wang and Guo, 2020] to motivate the direct estimator of the LOO AUROC,
@@
where $\mathcal{D}^{0}$ is the set of all negative observations and $\mathcal{D}^{1}$ is the set of all positive observations.
Weight stabilization: Often it is the case that using the computed posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$ as the proposal distribution for importance sampling has slow convergence properties – the $1/\ell$ importance weights are known to have large or unbounded variance [Peruggia, 1997], making the importance sampler estimate for LOO noisy.
Two practical model agnostic methods for controlling the tail of the importance weights are through weight truncation [Ionides, 2008] and Pareto smoothing  [Vehtari et al., 2024, 2017]. Pareto smoothing replaces the largest $M$ weights with their corresponding rank-values from a fitted generalized Pareto-distribution [Zhang and Stephens, 2009]. Pareto smoothed importance sample (PSIS)-based LOO implementations are widely available in software packages such as Stan and ArviZ. However, PSIS-LOO is insufficient in problems where the Pareto distribution does not well-fit the tail distribution of importance weights – where the estimated Pareto shape parameter $\hat{k}$ exceeds $0.7$ – this parameter is a useful criterion for accessing the validity of a given importance sampler. In such a case, it would be desirable to perform an additional model-specific controlled transformation on the proposal distribution that will induce more efficient computations.
@@ 2.1 Adaptive importance sampling
We apply the concept of adaptive importance sampling [Bugallo et al., 2017, Cornuet et al., 2011, Elvira and Martino, 2022] to transform the posterior distribution to be closer to the LOO distribution $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$ (relationships between the different distributions is depicted in Fig. 1).
Consider the bijection $T_{i}:\mathbb{R}^{p}\to\mathbb{R}^{p}$, defined for observation $i,$ and let $\boldsymbol{\phi}\equiv T_{i}(\boldsymbol{\theta}).$ By change of variables, $\pi_{\boldsymbol{\phi}}(\boldsymbol{\phi}|\ldots)=\pi\left(T_{i}^{-1}(\boldsymbol{\phi})|\ldots\right)\mathcal{J}_{i}^{-1}(\phi),$ where we denote $\mathbf{J}_{T}=\nabla T$, $\mathcal{J}_{T_{i}}^{-1}(\boldsymbol{\phi})=\left|\mathbf{J}_{T_{i}}^{-1}(\boldsymbol{\phi})\right|,$ and $\mathcal{J}_{T_{i}}(\boldsymbol{\theta})=\left|\mathbf{J}_{T_{i}}(\boldsymbol{\theta})\right|={1}/{\mathcal{J}_{T_{i}}^{-1}(\boldsymbol{\phi})},$ the determinants of the Jacobians of the inverse and forward transformations respectively. One can rewrite the expectation in Eq. 3 in terms of an integral over $\pi_{\boldsymbol{\phi}}$,
${\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})}{\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})}\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\theta}))}{\pi(T_{i}^{-1}(\boldsymbol{\theta})|\mathcal{D})}\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}.$ (9)
One can then define a Monte-Carlo approximation of Eq. 9 using importance sampling, by sampling $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$ so that $\boldsymbol{\phi}_{k}=T_{i}(\boldsymbol{\theta}_{k})\overset{\mathrm{iid}}{\sim}\pi_{\boldsymbol{\phi}}(\boldsymbol{\phi}|\mathcal{D}):$
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]$ $\approx\sum_{k=1}^{s}\frac{\eta_{ik}}{\sum_{j=1}^{s}\eta_{ij}}f(\boldsymbol{\phi}_{k})$ (10)
$\eta_{ik}$ $=\frac{\mathcal{J}_{T_{i}}(\boldsymbol{\theta}_{k})}{\ell(\boldsymbol{\phi}_{k}|\mathbf{x}_{i},y_{i})}\frac{\pi(\boldsymbol{\phi}_{k}|\mathcal{D})}{\pi(\boldsymbol{\theta}_{k}|\mathcal{D})}.$ (11)
@@
$\frac{\pi(\boldsymbol{\phi}|\mathcal{D})}{\pi(\boldsymbol{\theta}|\mathcal{D})}=\prod_{i}\frac{\ell(\boldsymbol{\phi}|\mathbf{x}_{i},y_{i})}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}.$ (12)
Computing this expression requires iterating over the entire dataset. For large datasets, one can turn to variational approximations.
Using variational posteriors: For computational expediency, variational methods are often used in place of MCMC for Bayesian inference, obtaining a variational approximation $\hat{\pi}(\boldsymbol{\theta}|\mathcal{D})$ to the true posterior, where $\hat{\pi}$ lies within a given family of transformed probability distributions. In problems where one expects a substantial discrepancy between the true posterior and $\hat{\pi},$ one may correct for this discrepancy by noting that
${\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})}{\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})}\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\theta}))}{\hat{\pi}(T_{i}^{-1}(\boldsymbol{\theta})|\mathcal{D})}\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$ (13)

### v3

# Perturbative partial moment matching and gradient-flow adaptive importance sampling transformations for Bayesian leave one out cross-validation

@@ Abstract
Importance sampling (IS) allows one to approximate leave one out (LOO) cross-validation for a Bayesian model, without refitting, by inverting the Bayesian update equation to subtract a given data point from a model posterior. For each data point, one computes expectations under the corresponding LOO posterior by weighted averaging over the full data posterior. This task sometimes requires weight stabilization in the form of adapting the posterior distribution via transformation. So long as one is successful in finding a suitable transformation, one avoids refitting. To this end, we motivate the use of bijective perturbative transformations of the form $T(\bm{\theta})=\bm{\theta}+hQ(\bm{\theta}),$ for $0<h\ll 1,$ and introduce two classes of such transformations: 1) partial moment matching and 2) gradient flow evolution. The former extends prior literature on moment-matching under the recognition that adaptation for LOO is a small perturbation on the full data posterior. The latter class of methods define transformations based on relaxing various statistical objectives: in our case the variance of the IS estimator and the KL divergence between the transformed distribution and the statistics of the LOO fold. Being model-specific, the gradient flow transformations require evaluating Jacobian determinants. While these quantities are generally readily available through auto-differentiation, we derive closed-form expressions in the case of logistic regression and shallow ReLU activated neural networks. We tested the methodology on an $n\ll p$ dataset that is known to produce unstable LOO IS weights.
@@ 1 Introduction
In Bayesian workflows, multiple models are often fitted to a given dataset, and a selection procedure is applied to decide which model will be the most consistent with future observations. Prediction accuracy is most naturally estimated using cross-validation (CV) of which many variants exist. Commonly, a model trained using a given partition of the available data and evaluated using the remaining unused data. However, estimates of out-of-sample model metrics using train-test splitting are statistically noisy (Dietterich, 1998; Kohavi, 1995) unless computationally expensive k-fold cross-validation (i.e., fitting the model multiple times on cross-over the entire dataset) is employed (Rodriguez et al., 2010; Wong & Yeh, 2020).
Although $N$-fold – also known as leave one out (LOO) – CV is the most expensive of $k$-fold estimators, there exist computationally efficient LOO techniques that completely avoid refitting. For example, the Akaike Information criteria (AIC) and Bayesian variants (Stone, 1977; Watanabe, 2010; Gelman et al., 2014; Watanabe, 2013) are asymptotic approximations of LOO CV. For Bayesian models, a more precise way to compute LOO CV is to use importance sampling (Vehtari et al., 2017; Piironen & Vehtari, 2017a), which works by using the full data posterior measure as a proposal distribution for each data point’s LOO posterior measure. However, in cases where the LOO measure and full measure are very different, importance sampling can fail (Piironen & Vehtari, 2017a). To ameliorate this possibility, we introduce an adaptive importance sampling methods for LOO CV based on using transformations that bring the proposal distribution closer to LOO posteriors under the principle that the transformation should be a small perturbation. We derive these transformations by defining gradient flows that minimize given statistical objective. While the transformations are model-dependent, the method is made completely general when using autograd for computing model gradients – for computational efficiency we derive the transformations exactly for a large class of classification models.
@@ 2 Preliminaries
2.1 Notation
We denote vectors (assumed to be column vectors unless otherwise stated) using bold-faced lowercase symbols, and matrices using bold-faced uppercase symbols. Given a matrix $\mathbf{W}=(w_{ij})$, the $i$-th row is denoted $\mathbf{w}_{i,}$ and $j$-th column is denoted $\mathbf{w}_{,j}$.
We refer to the entire set of observed training data as $\mathcal{D}=\{\mathbf{d}_{i}\}_{i=1}^{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. As shorthand, we denote the set of training data where the i-th observation is left out as $\mathcal{D}^{(-i)}=\mathcal{D}\setminus\{\mathbf{d}_{i}\}$. Expectations with respect to the posterior distribution of $\bm{\theta}$ are denoted $\mathbb{E}_{\bm{\theta}|\mathcal{D}}$, and with respect to the posterior distribution of $\bm{\theta}$ if observation $i$ is left out are denoted $\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}$.
@@ 2.1 Notation
Figure 1: Relationships between probability densities. One wants to sample from $\pi(\bm{\theta}|\mathcal{D}^{(-i)}),$ the LOO distribution for observation $i$, by sampling from the full-data posterior $\pi(\bm{\theta}|\mathcal{D}).$ The transformation $T_{i}$ on the full-data posterior brings the sampling distribution closer to the target LOO distribution.
2.2 Importance sampling-based approximate leave one out cross validation (IS-LOO)
@@
Suppose that one has pre-trained a Bayesian model such that one is able to sample its posterior parameters $\bm{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\bm{\theta}|\mathcal{D}).$ Our objective is to use knowledge of this full-data posterior distribution to estimate how the model would behave if any single point is left out at training. One can relate the full-data model to the model with observation $i$ left out using the Bayesian update equation
$\pi(\bm{\theta}|\mathcal{D})=\frac{\ell(\bm{\theta}|\mathbf{d}_{i})\pi(\bm{\theta}|\mathcal{D}^{(-i)})}{\int\ell(\bm{\theta}|\mathbf{d}_{i})\pi(\bm{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\bm{\theta}},$ (1)
@@
which is a Fredholm integral equation of the second-kind with respect to $\pi(\bm{\theta}|\mathcal{D}^{(-i)})$. This integral equation is in-practice difficult to solve due to the typically-high dimensionality of $\bm{\theta}.$ Note that Eq. 1 is equivalent to Bayes’ rule
$\pi(\bm{\theta}|\mathcal{D})\propto\pi(\bm{\theta})\prod_{i=1}^{n}\ell(\bm{\theta}|\mathbf{d}_{i}).$ (2)
@@
Rather than directly inverting Eq. 1 to obtain $\pi(\bm{\theta}|\mathcal{D}^{(-i)})$, our starting point is the observation that Eq. 1 implies
$\frac{\pi(\bm{\theta}|\mathcal{D}^{(-i)})}{\pi(\bm{\theta}|\mathcal{D})}=\frac{\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}[\ell(\bm{\theta}|\mathbf{d}_{i})]}{\ell(\bm{\theta}|\mathbf{d}_{i})}\equiv\nu_{i}(\bm{\theta}),$ (3)
@@
providing the ratio of densities functions between a distribution we know (the full-data posterior $\pi(\bm{\theta}|\mathcal{D})$) and a distribution whose statistics we would like to compute (the point-wise LOO posterior $\pi(\bm{\theta}|\mathcal{D}^{(-i)})$). To use the former to compute the latter we turn to Monte Carlo (Barbu & Zhu, 2020; Robert & Casella, 2013) – the use of statistical sampling to compute a desired quantity (typically an integral). Importance Sampling (IS) is a Monte Carlo method where one computes expectations with respect to a target distribution by taking a weighted average of samples with respect to a given proposal distribution. For an integrable function $f$,
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[f(\bm{\theta})\right]=\int f(\bm{\theta})\pi(\bm{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\bm{\theta}=\int f(\bm{\theta})\frac{\pi(\bm{\theta}|\mathcal{D}^{(-i)})}{\pi(\bm{\theta}|\mathcal{D})}\pi(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}=\mathbb{E}_{\bm{\theta}|\mathcal{D}}\left[f(\bm{\theta})\nu_{i}(\bm{\theta})\right].$ (4)
@@
We approximate Eq. 4 by sampling over $\bm{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\bm{\theta}|\mathcal{D})$, and computing the Monte Carlo integral
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}[f(\bm{\theta})]\approx\sum_{k=1}^{s}\nu_{ik}f(\bm{\theta}_{k})$ (5)
@@
$\nu_{ik}=\frac{\nu_{i}(\bm{\theta}_{k})}{\sum_{j=1}^{s}\nu_{i}(\bm{\theta}_{j})}=\frac{(\ell(\bm{\theta}_{k}|\mathbf{d}_{i}))^{-1}}{\sum_{k=1}^{s}(\ell(\bm{\theta}_{k}|\mathbf{d}_{i}))^{-1}},$ (6)
so that the undetermined constant $\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[\ell(\bm{\theta}|\mathbf{d}_{i})\right]$ cancels out. Eqs. 56 define a well-known (Gelfand et al., 1992) Monte Carlo estimator for LOO.
2.3 LOO cross validation based metrics
The Bayesian LOO information criterion (LOO-IC), of which the Aikaike Information Criterion (AIC) is an asymptotic approximation, can be computed via:
@@ 2.3 LOO cross validation based metrics
For classification problems, the out-of-sample area under the receiver operator curve or the precision-recall curve is often required. This can similarly be computed by propagating LOO estimates of the outcome probabilities
$\widehat{\mathscr{p}}_{\textrm{loo},i}=\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[\mathscr{p}_{i}(\bm{\theta})\right]\approx\sum_{k=1}^{s}{\nu}_{ik}\mathscr{p}(\bm{\theta}_{k},\mathbf{x}_{i}).$ (8)
2.4 Weight stabilization
Often it is the case that using the computed posterior $\pi(\bm{\theta}|\mathcal{D})$ as the proposal distribution for importance sampling has slow convergence properties – the $1/\ell$ importance weights, being fat tailed, are known to have large or unbounded variance (Peruggia, 1997), making the importance sampler estimate for LOO expectations (Eq. 6) noisy.
Two practical model agnostic methods for controlling the tail of importance weights are weight truncation (Ionides, 2008) and Pareto smoothing  (Vehtari et al., 2024; 2017). Pareto smoothing replaces the largest $M$ weights with their corresponding rank-values from a fitted generalized Pareto-distribution (Zhang & Stephens, 2009). Pareto smoothed importance sample (PSIS)-based LOO implementations are widely available in software packages such as Stan and ArviZ. However, PSIS-LOO fails when the tail distribution of importance weights is not well-fit by the Pareto distribution; a general rule of thumb is that the parameter $\hat{k}$ exceeds $0.7$. In these cases, performing an additional model-specific controlled transformation on the proposal distribution will induce more efficient computations. Later on, as in Paananen et al. (2021), we will use the estimated Pareto shape parameter $\hat{k}$ on post-transformation IS weights in order to evaluate the success of different transformations. Effective transformation should be able to reduce the Pareto shape parameter to below the given threshold.
@@ 2.5 Adaptive importance sampling
We apply the concept of adaptive importance sampling (Bugallo et al., 2017; Cornuet et al., 2011; Elvira & Martino, 2022) to transform the posterior distribution to be closer to the LOO distribution $\pi(\bm{\theta}|\mathcal{D}^{(-i)})$ (the relationships between the different distributions are depicted in Fig. 1).
Consider the bijection $T_{i}:\mathbb{R}^{p}\to\mathbb{R}^{p}$, defined for observation $i,$ and let $\bm{\phi}\equiv T_{i}(\bm{\theta}).$ By change of variables, $\pi_{\bm{\phi}}(\bm{\phi}|\ldots)=\pi\left(T_{i}^{-1}(\bm{\phi})|\ldots\right)\mathcal{J}_{i}^{-1}(\phi),$ where we denote $\mathbf{J}_{T}=\nabla T$, $\mathcal{J}_{T_{i}}^{-1}(\bm{\phi})=\left|\mathbf{J}_{T_{i}}^{-1}(\bm{\phi})\right|,$ and $\mathcal{J}_{T_{i}}(\bm{\theta})=\left|\mathbf{J}_{T_{i}}(\bm{\theta})\right|={1}/{\mathcal{J}_{T_{i}}^{-1}(\bm{\phi})},$ The expectation in Eq. 4 in terms of an integral over $\pi_{\bm{\phi}}$ is
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[f(\bm{\theta})\right]$ $=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\pi(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\frac{\pi(\bm{\theta}|\mathcal{D})}{\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})}\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}$
$=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\frac{\pi(\bm{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\bm{\theta}))}{\pi(T_{i}^{-1}(\bm{\theta})|\mathcal{D})}\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}.$ (9)
Define a Monte Carlo approximation of Eq. 9 using importance sampling, by sampling $\bm{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\bm{\theta}|\mathcal{D})$ so that $\bm{\phi}_{k}=T_{i}(\bm{\theta}_{k})\overset{\mathrm{iid}}{\sim}\pi_{\bm{\phi}}(\bm{\phi}|\mathcal{D}):$
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[f(\bm{\theta})\right]\approx\sum_{k=1}^{s}\frac{\eta_{ik}}{\sum_{j=1}^{s}\eta_{ij}}f(\bm{\phi}_{k})\qquad\eta_{ik}=\frac{\mathcal{J}_{T_{i}}(\bm{\theta}_{k})}{\ell(\bm{\phi}_{k}|\mathbf{d}_{i})}\frac{\pi(\bm{\phi}_{k}|\mathcal{D})}{\pi(\bm{\theta}_{k}|\mathcal{D})}.$ (10)
@@
$\frac{\pi(\bm{\phi}|\mathcal{D})}{\pi(\bm{\theta}|\mathcal{D})}=\frac{\pi(\bm{\phi})}{\pi(\bm{\theta})}\prod_{i}\frac{\ell(\bm{\phi}|\mathbf{d}_{i})}{\ell(\bm{\theta}|\mathbf{d}_{i})}.$ (11)
Computing this expression requires iterating over the entire dataset. There are various methods to avoid this expensive computation, for instance also using Monte Carlo approximation by sampling data points. For large datasets, one can turn to variational approximations.
2.6 Correcting variational posteriors
For computational expediency, variational methods are often used in place of MCMC for Bayesian inference, obtaining a variational approximation $\hat{\pi}(\bm{\theta}|\mathcal{D})$ to the true posterior, where $\hat{\pi}$ lies within a given family of probability distributions. In problems where one expects a substantial discrepancy between the true posterior and $\hat{\pi},$ one may correct for this discrepancy by noting that
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[f(\bm{\theta})\right]$ $=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\pi(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\frac{\pi(\bm{\theta}|\mathcal{D})}{\hat{\pi}_{\bm{\phi}}(\bm{\theta}|\mathcal{D})}\hat{\pi}_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}$
$=\int f(\bm{\theta})\nu_{i}(\bm{\theta})\frac{\pi(\bm{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\bm{\theta}))}{\hat{\pi}(T_{i}^{-1}(\bm{\theta})|\mathcal{D})}\hat{\pi}_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\mathrm{d}\bm{\theta}$ (12)